please complete each of the following bombardment reactions.

on + 338 U →? Express your answer as an isotope or a particle. 19 ADD A = O O ?
η +? - ??Na + 3He Express your answer as an isotope or a particle. - ΑΣΦ 6 - Ο ΕΞ ?
Answer #1

1)

Let the product be X

Balance mass number on both sides

mass of n + mass of U = mass of X

1*1 + 1*238 = M

M = 239

Balance atomic number on both sides

atomic of n + atomic of U = atomic of X

1*0 + 1*92 = A

A = 92

Atomic number 92 is for element U

So, the product is

239-U

Answer: 23992U

2)

Let the reactant be X

Balance mass number on both sides

mass of n + mass of X = mass of Na + mass of He

1*1 + M = 1*24 + 1*4

M = 27

Balance atomic number on both sides

atomic of n + atomic of X = atomic of Na + atomic of He

1*0 + A = 1*11 + 1*2

A = 13

Atomic number 13 is for element Al

So, the reactant is

27-Al

Answer: 2713Al